ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

13. Combinatorics

1991 MS Classification: 1991 AMS Classification: 05 Khosrovshahi, Gholamreza B. , Institute for Studies in Theoreretical Physics and Mathematics (IPM), Tehran, Iran On trades and their codes We consider the families of t-subsets vesus k-subsets inclusion matrices. We call the null spaces of these matrices over any finite field trade spaces. Trades have some nice combinatorial properties. Some of these spaces generate codes with some nice properties. These codes are a generalization of graphical codes which were recently studied very closely. In this talk we examine some examples of such codes.